OpenStudy (anonymous):

im lost ... −16t2+80x−60=0 Solving for t, t=(5±10−−√)2

OpenStudy (lgbasallote):

why do you have mixed variables -__-

OpenStudy (campbell_st):

well simplify the problem -4(4t^2 - 20t +15)

OpenStudy (anonymous):

Ah this is the a nswer i gave you earlier for the height?

OpenStudy (anonymous):

yeah im lost

OpenStudy (campbell_st):

so using the GQF \[t = (20\pm \sqrt{(-20)^2-4\times4\times15})/(2 \times4)\] \[t = (20 \pm \sqrt{16 \times10})/8 = (5\pm \sqrt{10})/2\]

OpenStudy (anonymous):

umm yeah im already there

OpenStudy (anonymous):

Screw it the equation th ing is broken

OpenStudy (anonymous):

Use the quadratic formula to s olve for t

OpenStudy (anonymous):

\[t=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

I dont think ive learned that yet

OpenStudy (anonymous):

in my class

OpenStudy (anonymous):

Is this a physics problem or calculus?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

You should've learned the quadratic formula back in algebra

OpenStudy (anonymous):

woops....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so i would just plug in the numbers that were found in the early part of the question?

OpenStudy (anonymous):

\[ax^2+bx+c=0\], you can only use the quadratic formula when you have your equation like this

OpenStudy (anonymous):

in this case you do, and you can also divide out a 4 from all the terms 16,80, and 60

OpenStudy (anonymous):

\[-4t^2+20x-15=0\] a=-4 b=20 c=-15

OpenStudy (anonymous):

ok? so now I can solve by getting

OpenStudy (anonymous):

by itself

OpenStudy (anonymous):

T

OpenStudy (anonymous):

no, use the quadratic formula as it is written above right now, \[t=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[t=\frac{-20 \pm \sqrt{20^2-4(-4)(-15)}}{2(-4)}\]

OpenStudy (anonymous):

\[t=\frac{-20 \pm \sqrt{160}}{-8}=\frac{-20 \pm 4\sqrt{10}}{-8}=\frac{-5 \pm \sqrt{10}}{-2}=\frac{5 \pm \sqrt{10}}{2}\]

OpenStudy (anonymous):

t= -20 sqrt 360/-8

OpenStudy (anonymous):

look above

OpenStudy (anonymous):

wow im way off!!

OpenStudy (anonymous):

Remember to keep the positive and negative signs for a,b,c

OpenStudy (anonymous):

so now I have 3.16227766/2 = 1.58113883

OpenStudy (anonymous):

5+ or - = 1.58113883

OpenStudy (anonymous):

\[t \approx .9189s, t \approx 4.081s\]

OpenStudy (anonymous):

can you give me the all the numbers shown on your calculator

OpenStudy (anonymous):

They are irrational numbers, there are infinitely many decimal places

OpenStudy (anonymous):

Use the answer with the square root

OpenStudy (anonymous):

That is accurate

OpenStudy (anonymous):

no just the numbers shown on your calculator bcause I need the precise number

OpenStudy (anonymous):

You can't get a precise number from an irrational number! Use \[t=\frac{5 \pm \sqrt{10}}{2} s\]

OpenStudy (anonymous):

That is the precise answer, any decimal answer is an approximation

OpenStudy (anonymous):

what I mean is when plug the answers into the my computer it needs the closets thing to an approximate number

OpenStudy (anonymous):

so it must have the all the numbers shown from the calculator for it to be correct

OpenStudy (anonymous):

Your calculator can show hundreds of digits if you want it to

OpenStudy (anonymous):

How many decimal places do you need?

OpenStudy (anonymous):

\[t \approx 0.918861169916s, t \approx 4.08113883008s\]

OpenStudy (anonymous):

like im using a standard calcultor so it only shows like 9 or 10 numbers

OpenStudy (anonymous):

here are my instructions : For example, the fraction 2/3 is exact, while the decimals 0.67, 0.666666667 and 0.666666666666666666667 are all approximate values for 2/3. Where an exact answer is called for, an approximate answer will be marked wrong.

OpenStudy (anonymous):

So if you use a decimal here, you will be wrong

OpenStudy (anonymous):

yup lol

OpenStudy (anonymous):

Use the answer i gave you with the square root

OpenStudy (anonymous):

so like t t= .9189 just sqrt it

OpenStudy (anonymous):

...

OpenStudy (anonymous):

\[t=\frac{5+\sqrt{10}}{2}, t=\frac{5-\sqrt{10}}{2}\]

OpenStudy (anonymous):

Those are the exact answers for the time

OpenStudy (anonymous):

ok got it , thanks sorry for the confusion !!